OpenStudy (anonymous):

Solve A = 5(b + 2c) for c in terms of A and b.

OpenStudy (anonymous):

lots of ways to do this we can divide by 5 to get \[\frac{A}{5}=b+2c\] subtract b to get \[\frac{A}{5}-b=2c\] divide by 2 to get \[\frac{A}{10}-\frac{b}{2}=c\] that is one way

OpenStudy (anonymous):

A=5b+10c c=(A-5b)/10

OpenStudy (anonymous):

or we can multiply out first and get \[A=5B+10c\] then subtract \(5b\) and get \[A-5b=10c\] divide by 10 and get \[\frac{A-5b}{10}=c\]

OpenStudy (anonymous):

whatever you like best